lilbaddieeeeee:

A local newspaper claims that the number of flu cases is increasing exponentially. On Monday, there were 8 flu cases reported. On Tuesday, there were 12 flu cases and on Wednesday there were 24 cases reported. On Thursday, there were 30 flu cases. Was the newspaper's claim of exponential increase accurate?

lilbaddieeeeee:

plz help

jhonyy9:

Monday was 2*4 = 2*2^2 Tuesday 12 = 3*4 = 3*2^2 Wednesday 24 = 2*3*2^2 Thursday 30 = 2*3*5 any idea ?

jhonyy9:

@Laylalyssa